OpenStudy (anonymous):

Tom plans to mow a lawn today to earn some extra money. She estimates that she will spend $1.80 on fuel and $3.45 on lunch. She will charge $11.50 to mow the lawn. What will be her profit for mowing the lawn? A. $5.25 B. $6.25 C. $9.85 D. $16.75

OpenStudy (anonymous):

someone please help me

OpenStudy (carolina4567):

B. 6.25

OpenStudy (carolina4567):

Tom plans to mow a lawn today to earn some extra money. She estimates that she will spend $1.80 on fuel and $3. 45 on lunch. She will charge 11.50 to mow the lawn. What will be her profit for mowing the lawn? Profit = $11.50 -( $1.80 + $3.45) = 11.50 - 5.25 = $6.25

OpenStudy (anonymous):

thank you so much (:

OpenStudy (carolina4567):

YW :)